1 / (e^x + 1) What is the solution? I also want to know somethng related to solving integration... When I look at a problem, i dunno how to solve it... bu wen i look at the solution steps then i m able to understand the steps. wat is the problem that i shud overcome so that i can solve problems?

1 / (e^x + 1)
What is the solution? I also want to know somethng related to solving integration... When I look at a problem, i dunno how to solve it... bu wen i look at the solution steps then i m able to understand the steps. wat is the problem that i shud overcome so that i can solve problems?

Grade:12

1 Answers

AskiitiansExpert Abhinav Batra
13 years ago

Dear Mohini

let e^x + 1=t

then e^x dx=dt

so dx/e^x+1=dt/e^x*t=dt/(t-1)t=dt/t-1-dt/t

∫dt/t-1-dt/t=ln(t-1/t)+c=ln(e^x/e^x + 1)+c

Good Luck

AskIITIANS Expert

Abhinav Batra